OpenStudy (anonymous):

The guy wire of a telephone pole meets the pole at a height of 5 meters and connects to the ground 4 meters from the pole. Assuming that the ground is flat, what is the height of the guy wire halfway between the pole and the point at which it meets the ground?

OpenStudy (campbell_st):

draw a picture  there seem to be 3 parts to the problem. 1. Find the angle between the guy rope and ground using trig \[\tan(\theta) = \frac{5}{4} \] 2. find the length of the guy wire using trig \[y^2 = 5^2 + 4^2\] 3. Find the height of x after finding 1/2 of the guy wire using \[\sin(\theta) = \frac{x}{1/2\times hypotenuse}\]